Suppose that the engine of a car delivers a maximum force of 14,000 N. 1) In the absence of any other forces, what is the maximum acceleration this engine can produce in a 1250-kg car? (Express your answer to two significant figures.) m/s? Submit copyright 2017 Freeman publishers a division of an Learning. About Worth I Tech support I Find Your Local Sales Rep Terms of Use I Privacy policy
media%2F717%2F717eeaa9-0e3f-43a9-83c5-d3
Answer #1

According to the question we need to find the value of the astronaut weight on mars.

mass = 77 kg g = 0.38 * 9.8 m/s^2 = 3.724 m/s^2

Force = mass * g = 77 kg * 3.724 m/s^2 = 286.748 Newton ==================ANSWER()_